OpenStudy (anonymous):

Solve the equation will give medal! 3x-11=-18

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

that's not a answer choice for mine

OpenStudy (anonymous):

The point is for you to solve it, and us to help you if you get stuck. What would you do first? Or, what answer do you get? Code of conduct clearly says: "OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. Have homework problems? We can help you - as long as you are looking to learn how to get your solution (and not just the answer)."

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

my answer choices are A- x=-17 B- x=-5 C- x=-3 D- x=5

OpenStudy (anonymous):

1) Again, have you tried doing it on your own, what do you get, where do you have problems? 2) Unless there's some missing information or something was transcribed wrong none of those answers are correct.